OpenStudy (anonymous):

If A=10 and given the expression is A=A+1, then the next value of A will be?

OpenStudy (anonymous):

It's either 10 or 11 it confuses me XD

OpenStudy (anonymous):

A=10+1 A=11

OpenStudy (anonymous):

But wouldn't it be 10=10+1?

OpenStudy (anonymous):

The next value of A is 9.

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

A=11

OpenStudy (anonymous):

10 = 11

OpenStudy (anonymous):

I think @Sargams14 is right.

OpenStudy (anonymous):

oh it is